3(a) A machine has a life of 20 years, costs $200,000 and has an estimated salvage value of $10,000. (i) For the Straight Line method of depreciation, what is the depreciation rate and what is the book value at the end of year 10? (ii) If the declining balance method is to be used, at what depreciation rate (i.e. the capital cost allowance CCA rate), the book value at the end of year 10 will be the same as for the straight-line method?
Answer #1

Answer:

A)

Straight line depreciation rate=1/useful life=1/20=5%

Depreciable base=200000-10000=$190000

Depreciation per year=190000*5%=$9500

Book Value at the end of 10 year=200000-9500*10=$105000

B)

let r be the rate of depreciation

Asset cost A=200000

Then book vale at the end of 10 year under declining balance method= A*(1-r)^10

Since

A*(1-r)^10=105000

200000*(1-r)^10=105000

r=6.24%